OpenStudy (anonymous):

x^2 - 3x = -8 please solve using the quadratic formula

OpenStudy (mimi_x3):

\[x = \frac{-(-3)\pm\sqrt{(-3)^{2}-4(1)(8)}}{2(1)} = \]

OpenStudy (ash2326):

x^2-3x+8=0 ax^2+bx+c= standard form of a quadratic equation \[x=(-b \pm \sqrt (b^2-4ac))/2a\] a=1 b=-3 c=8 \[x=(3 \pm \sqrt (9-4*8))/2\] \[x=(3 \pm \sqrt (-23))/2\] \[x=(3 \pm \sqrt 23 i)/2\]